OpenStudy (anonymous):

PLEASE HELP ALGEBRA 2

OpenStudy (anonymous):

rationalize the denominator and simplify \[\frac{ \sqrt{10} }{ \sqrt{3}}\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

lease multiply both numerator and denminator by sqrt(3), namely: \[\frac{ \sqrt{10} }{ \sqrt{3} }=\frac{ \sqrt{10} *\sqrt{3}}{ \sqrt{3}*\sqrt{3} }=...\]

OpenStudy (anonymous):

so \[\frac{ \sqrt{30} }{ 1}\]

OpenStudy (michele_laino):

sorry, wait please:

OpenStudy (michele_laino):

\[=\frac{ \sqrt{30} }{ 3 }\] since\[\sqrt{3}*\sqrt{3}=3\]

OpenStudy (anonymous):

is that it?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

\[\frac{ \sqrt{30} }{ 3 }\]